OpenStudy (anonymous):

what is the sum of 5x and 9x?

OpenStudy (anonymous):

5x + 9x = 14x

OpenStudy (anonymous):

\(\Large 5+9=14\) \(\Large \Rightarrow 5x + 9x = 14x \)